# Single Image Rolling Shutter Removal with Diffusion Models

### v1

@@ III-D Rolling Shutter Removal Module
IV Experiment
This section is organized as follows: We begin by the implementation details in Section IV-A and outlining both existing and our proposed datasets in Section IV-B. We then compare our method’s performance against other approaches on various benchmarks in Section IV-C, with qualitative evaluations presented in Section IV-D. The performance of our model with global shutter images as input was evaluated in Section IV-E. The generalizability of our approach is examined through experiments across different devices in Section IV-F. An in-depth analysis of our framework’s intricate design features is provided in Section IV-G.
IV-A Implementation Details
Our model is implemented in PyTorch and trained on 8 NVIDIA 2080Ti GPUs. During the training phase, we utilized the Adam optimizer with a learning rate of $3\times 10^{-4}$ and a batch size of 64. The backbone diffusion model settings include an input resolution of $64\times 64$, 1,000 diffusion steps, a linear noise scheduler, and the target of predicting “$x_{0}$” as the generating object. The entire training process comprised 150,000 steps and took approximately 40 hours. In the forward diffusion process, we performed 8 sampling steps with $\eta=0$. The inference time can be up to 28.1 ms per frame on a single NVIDIA 2080Ti.
IV-B Dataset
In our experiments, we validate the proposed RS correction method on 3 datasets, including one long-standing public datasets and two new ones for comprehensive evaluation. To ascertain the reliability of our approach, we synchronize our evaluations with established RS image correction techniques [7, 10, 16], utilizing the Building dataset [71, 72, 73] as a standard benchmark. Additionally, we extend our test scope with a challenging dataset RS-Homo from Yan et al. [16], crafted to simulate adverse conditions like low lighting and scant textures, critical for testing diffusion model robustness. A notable contribution of this paper is the RS-Real dataset, detailed in Section III-B, which exploits sensor information during image capture to address common synchronization challenges and offers a new methodology for data gathering in RS scenarios.
IV-C Quantitative Comparison
In our study, we evaluate our RS image correction method using key metrics for visual quality and motion accuracy. Visual comparisons are done via the Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) against GT GS images. The motion-based assessment considers the Endpoint Error (EPE) for correction flow accuracy. Pixels on black edges, lacking information, are omitted from our analysis. Notably, diffusion models can produce different results with each inference. Therefore, we report both the mean and the standard deviation of the results over 10 inference iterations.
TABLE I: Comparison of PSNR, SSIM, and EPE between our method and existing ones on Building dataset. The values enclosed within parentheses indicate the standard deviation of our results over 10 inference iterations.
@@ IV-C Quantitative Comparison
Rengarajan et al. [6] 29.82 0.67 11.89
Purkait et al. [8] 29.22 0.55 8.32
Grundmann et al. [74] 32.57 0.72 3.34
Rengarajan et al. [7] 32.25 0.70 3.76
Kandula et al. [10] 32.85 0.73 2.84
Yan et al. [16] 33.34 0.75 1.25
RS-Diffusion 34.92(±$\pm 0.045$) 0.79(±$\pm 0.002$) 0.90(±$\pm 0.018$)
Building dataset. Our study compares with traditional and deep learning-based RS correction methods, as shown in Table I. Approaches using curve detection [6, 8] to discern motion patterns struggle in low-structure environments, showing subpar PSNR and SSIM, alongside elevated EPE. The homography mixtures method [74] fares better with its video sequence foundation and refined feature correspondences. However, learning methods [7, 10] have proven robust to strong outliers and complex camera motions. Yan et al. model [16], merging homography mixtures with learning, showcases significant strides in metric performance due to constrained motion spaces from learned bases. Our proposed technique outperforms existing methods, setting a new state-of-the-art (SOTA) across evaluated metrics and emphasizing its effectiveness.
RS-Homo dataset. The results in Table II show our diffusion models-based framework substantially improves upon Yan et al.’s state-of-the-art single-image model. Demonstrating great proficiency in managing challenging scenarios, our proposed method surpasses specifically designed architectures, enhancing the capability of diffusion models to address rolling shutter (RS) effects effectively.
Figure 5: Comparison with existing methods: Rengarajan et al. [7], Rengarajan et al. [6], Purkait et al. [8], Grundmann et al. [74] and Kandula et al. [10] on a RS building image. Red vertical lines to highlight correction results.
Figure 6: Comparison on RS-Homo dataset. Column 1 shows the input RS image and Column 2 shows the ground-truth GS image. Column 3 and 5 are results by Yan et al. [16] and our method with the alignment heatmaps that use darker shades to indicate greater similarity compared to ground-truth GS images.
Figure 7: Comparison on RS-Real dataset. Column 1 shows the input RS image and Column 2 shows the ground-truth GS image. Column 3 and 5 are results by Yan et al. [16] and our method with the alignment heatmaps that use darker shades to indicate greater similarity compared to ground-truth GS images.
TABLE II: Comparison of PSNR, SSIM, and EPE between our method and Yan et al. [16] on the RS-Homo dataset. The values enclosed within parentheses indicate the standard deviation of our results over 10 inference iterations.
@@
Yan et al. [16] 26.15 0.77 4.10
RS-Diffusion 36.66(±$\pm 0.092$) 0.94(±$\pm 0.002$) 1.02(±$\pm 0.083$)
TABLE III: Comparison of PSNR, SSIM, and EPE between our method and Yan et al. [16] on the RS-Real dataset. The values enclosed within parentheses indicate the standard deviation of our results over 10 inference iterations.
@@
Yan et al. [16] 18.48 0.55 4.18
RS-Diffusion 22.02(±$\pm 0.031$) 0.70(±$\pm 0.001$) 2.12(±$\pm 0.010$)
@@
RS-Real dataset. Our dataset, designed for realism in content, RS-motion, and label accuracy, comprises 44,000 training and 1,000 test pairs across diverse scenes. It includes RS and GS images along with GT flow. We benchmark against Yan et al. [16] by retraining their model using the default settings on our trainset. The results confirm that our approach consistently outperforms theirs in both photometric and motion measures as demonstrated in Table III.
IV-D Qualitative Comparison
@@
Our method is evaluated against current methods on different benchmarks. We follow the settings used by Yan et al. [16] to compare with existing methods [7, 6, 8, 74, 10] in Fig. 5, and ours outperforms the others. In addition, due to limited open-source options, we compare with Yan et al. method on the remaining datasets. We visualize RS and ground-truth GS images in the first two columns, then present Yan et al. results paired with alignment heatmaps that use darker shades to indicate greater similarity compared to GT images. We then display our own results alongside their corresponding alignment heatmaps, illustrating our approach’s consistency with benchmarks and precision in alignment. This visual comparison highlights the effectiveness of our methodology.
RS-Homo Dataset. We select seven representative examples from the dataset in Fig. 6. In the first three rows, we display RS images from a typical scenario characterized by rich textures and evident RS motion. In this case, our method consistently outperforms theirs. In the fourth row, we present an RS image captured in a scene with moving objects. Our method demonstrates its capability to handle such challenging conditions, delivering superior results as evidenced by the RS-corrected images and their alignment heatmaps. In the fifth to seventh rows, we present examples from low-texture scenes and nighttime scenes. Notably, our method excels at producing accurate corrections, in contrast to the comparison method, which tends to over-correct the RS image.
RS-Real Dataset. The new benchmark comprises more challenging RS images and varied RS motions as shown in Fig. 7. The first four rows compare performances for different RS directions; right-skewed and left-skewed. Our method outperforms Yan et al. [16], proving its effectiveness. The fifth and sixth rows feature RS images from quick device movement that produces intricate motion patterns, such as curved buildings. Our method effectively corrects the RS motion while the comparison method fails. The final three rows present complex scenarios with moving cars; the comparative method struggles with this complexity, resulting in incomplete corrections, whereas our framework efficiently manages these dynamic elements. This selection highlights our approach’s robustness across diverse and realistic RS challenges.
IV-E Experiment of testing on global shutter images
In previous experiments, our model was trained on RS-IGF-GS pairs, thus naturally performing well in converting images from rolling shutter to global shutter. Additionally, we are curious about the performance of the trained model when given global shutter images as input. Therefore, we conducted experiments, the results of which are presented in the second row of Table IV and the second column of Fig. 8. Although the experimental results demonstrate that our model does not easily handle GS images, we believe the main reason lies in the absence of global shutter images during training.
To address this issue, we randomly sample 20% of the training data, utilizing the GS frames. The corresponding intra-gyro-fields are set to be all-zero, and we fine-tune our model for an additional 50,000 steps. More specifically, we crop 300$\times$400 patches from the 600$\times$800 GS images. The experimental results indicate that, after fine-tuning, our model gains the ability to handle global shutter inputs effectively, as illustrated in the last column of Fig. 8 and the third row of Table IV. Furthermore, the ability of the model to correct RS frames is further enhanced.
Figure 8: Comparison illustration of global shutter inputs before and after fine-tuning. The first row represents the input global shutter images, the second row depicts the correction effect before model fine-tuning, and the third row shows the correction effect after model fine-tuning.
IV-F Generalizability Experiments
Figure 9: The illustration of results that we leverage the pre-trained model on RS-Real to zero-shot inference on newly captured RS images using other mobile devices.
Experiments involve capturing new RS images with a “ONEPLUS 11” smartphone. These images are subsequently subjected to zero-shot inference using a model that has been pre-trained on the RS-Real dataset. It is important to note that the RS-Real dataset comprises images captured using a “XIAOMI CC9Pro” smartphone. The outcomes of these experiments are illustrated in Fig. 9.
Despite the fact that RS artifacts often correlate with specific sensors and devices, our framework’s robust backbone architecture ensures its generalizability across different mobile devices. This generalizability is critical for practical applications, as it demonstrates that the model can effectively handle RS images captured from a variety of devices without significant loss of performance. This capability is particularly valuable in real-world scenarios where the diversity of capturing devices can pose a significant challenge. By leveraging the robust features learned from the RS-Real dataset, our model can maintain high accuracy and reliability, regardless of the source device used for image capture.
TABLE IV: Comparison of PSNR, SSIM, and EPE between our model before Fine-tuning and after Fine-tuning on the RS-Real dataset. The input data types include rolling shutter and global shutter. The values enclosed within parentheses indicate the standard deviation of our results over 10 inference iterations.
[…]
Before Fine-tuning RS 22.02(±$\pm 0.031$) 0.70(±$\pm 0.001$) 2.12(±$\pm 0.010$)
GS 16.78(±$\pm 0.528$) 0.47(±$\pm 0.532$) 7.56(±$\pm 0.425$)
After Fine-tuning RS 22.25(±$\pm 0.023$) 0.71(±$\pm 0.001$) 2.06(±$\pm 0.017$)
GS 38.65(±$\pm 0.025$) 0.98(±$\pm 0.031$) 0.097(±$\pm 0.005$)
IV-G Ablation Studies
@@
We evaluate our framework design through experiments, starting with comparisons under diverse generative objectives, paired with analytical analyses. The effectiveness of the Patch-Attention Block is then scrutinized. Lastly, we explore the trade-off between performance and resolution.
IV-G1 Predicting IGF vs. GS Image
We contrast our “image-to-motion” pipeline with traditional “image-to-image” methods that use diffusion models to convert RS images to GS images at a fixed resolution of 256$\times$256, later upscaled to 600$\times$800 for visual metric comparison with GT GS images, as shown in Table. V. Our approach significantly outperforms the conventional framework in the RS removal task for two main reasons: 1) It is not constrained by diffusion model resolutions, since we can upsample IGF to correct RS images at their original size; 2) While “image-to-image” models learn the joint probability distribution between RS and GS images, our method gains additional improvements by learning from the distribution involving RS images, GS images, and IGFs, thus leveraging more information for enhanced results.
TABLE V: Comparison between different generating objects. “image-to-image” indicates transforming RS images to GS ones. “image-to-motion” represents predicting IGFs from RS images.
@@
IV-G2 Rolling-Shutter Patch Attention
As illustrated in Table. VI, our proposed Patch Attention module is effective as it not only aggregates features within patches, but also interacts with features between patches. The experimental results are also consistent with our a priori knowledge of RS motion.
TABLE VI: The effectiveness of Patch Attention. Intra refers to intra-patch attention and Inter indicates inter-patch attention.
@@ IV-G2 Rolling-Shutter Patch Attention
$\checkmark$ $\checkmark$ 22.02 0.70 2.12
IV-G3 Different Resolutions
Experimenting with different resolutions, our findings are summarized in Table. VII. Notably, the benefit of higher resolution plateaus at 64, with any further increase leading to a decrease in results. This can be attributed to the global nature of RS motion, which only necessitates sufficient resolution rather than very high definition.
TABLE VII: Comparison between data resolution and performance.
[…]
32 21.30 0.66 2.37
64 22.02 0.70 2.12
128 21.37 0.66 2.36
V Limitation
In the proposed RS-Real dataset, rolling shutter effects primarily result from camera rotation, though many instances also stem from camera translation and moving objects. Since the gyroscope records only the camera’s rotation, our dataset concentrates on rotation-induced RS images. It neglects the rolling shutter effects and corresponding correction flows caused by camera translation and moving objects, representing this work’s major limitation. Despite these limitations in our dataset and model, we believe that our dataset still provides an effective benchmark for the single-frame rolling shutter correction task, capable of advancing the field.
VI Conclusion
In this work, we have presented RS-Diffusion, the first Diffusion Model based approach for single frame rolling shutter rectification. We have captured a novel dataset, namely RS-Real, specifically tailored for this task. The RS-Real dataset contains captured RS images, and the corresponding ground-truth GS images can be created according to synchronized gyroscope data that recorded during the RS frame capturing, yielding RS-GS image pairs. Uniquely, this dataset fulfills both accuracy and authenticity requirements for RS research. In addition, we have presented the first diffusion models based framework for real-time RS correction using just one RS frame. We have achieved state-of-the-art performances when compared to previous single RS correction methods. To facilitate further research, we will publicly share our code and dataset with the community.

### v2

@@ Introduction
Experiments
We begin by outlining existing and our proposed datasets in Section Dataset. Our method’s performance is compared against others on varied benchmarks in Section Quantitative Comparison, with qualitative evaluations presented in Section Qualitative Comparison. An in-depth analysis of our framework’s intricate designs is described in Section Ablation Studies. Finally, we list implementation details and demonstrate the utility of our framework in a video stabilization application (Liu et al. 2013) within the supplementary materials.
Dataset
In our experiments, we validate the proposed RS correction method on 3 datasets, including one long-standing public datasets and two new ones for comprehensive evaluation. To ascertain the reliability of our approach, we synchronize our evaluations with established RS image correction techniques (Rengarajan, Balaji, and Rajagopalan 2017; Kandula, Kumar, and Rajagopalan 2020; Yan et al. 2023), utilizing the Building dataset (Xiao et al. 2010; Shao, Svoboda, and Van Gool 2003; Philbin et al. 2007) as a standard benchmark. Additionally, we extend our test scope with a challenging dataset RS-Homo from Yan et al. (Yan et al. 2023), crafted to simulate adverse conditions like low lighting and scant textures, critical for testing diffusion model robustness. A notable contribution of this paper is the RS-Real dataset, detailed in Section Intra Gyro Field, which exploits sensor information during image capture to address common synchronization challenges and offers a new methodology for data gathering in RS scenarios.
Quantitative Comparison
In our study, we evaluate our RS image correction method using key metrics for visual quality and motion accuracy. Visual comparisons are done via the Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) against GT GS images. The motion-based assessment considers the Endpoint Error (EPE) for correction flow accuracy. Pixels on black edges, lacking information, are omitted from our analysis.
Figure 5: Comparison with existing methods: 1. Rengarajan et al. (Rengarajan, Rajagopalan, and Aravind 2016), 2. Rengarajan et al. (Rengarajan, Balaji, and Rajagopalan 2017), 3. Purkait et al. (Purkait, Zach, and Leonardis 2017), 4. Grundmann et al. (Grundmann et al. 2012) and 5. Kandula et al. (Kandula, Kumar, and Rajagopalan 2020) on a RS building image. Red vertical lines to highlight correction results.
@@
1. Rengarajan et al. 29.82 0.67 11.89
2. Purkait et al. 29.22 0.55 8.32
3. Grundmann et al. 32.57 0.72 3.34
4. Rengarajan et al. 32.25 0.70 3.76
5. Kandula et al. 32.85 0.73 2.84
6. Yan et al. 33.34 0.75 1.25
RS-Diffusion 34.92 0.79 0.90
Table 1: Comparison of PSNR, SSIM, and EPE between our method and existing ones: 1. Rengarajan et al. (Rengarajan, Rajagopalan, and Aravind 2016), 2. Purkait et al. (Purkait, Zach, and Leonardis 2017), 3. Grundmann et al. (Grundmann et al. 2012), 4. Rengarajan et al. (Rengarajan, Balaji, and Rajagopalan 2017), 5. Kandula et al. (Kandula, Kumar, and Rajagopalan 2020) and 6. Yan et al. (Yan et al. 2023) on Building dataset.
Building dataset. Our study compares with traditional and deep learning-based RS correction methods, as shown in Table 1. Approaches using curve detection (Rengarajan, Rajagopalan, and Aravind 2016; Purkait, Zach, and Leonardis 2017) to discern motion patterns struggle in low-structure environments, showing subpar PSNR and SSIM, alongside elevated EPE. The homography mixtures method (Grundmann et al. 2012) fares better with its video sequence foundation and refined feature correspondences. However, learning methods (Rengarajan, Balaji, and Rajagopalan 2017; Kandula, Kumar, and Rajagopalan 2020) have proven robust to strong outliers and complex camera motions. Yan et al. model (Yan et al. 2023), merging homography mixtures with learning, showcases significant strides in metric performance due to constrained motion spaces from learned bases. Our proposed technique outperforms existing methods, setting a new state-of-the-art (SOTA) across evaluated metrics and emphasizing its effectiveness.
@@
Yan et al. 26.15 0.77 4.10
RS-Diffusion 36.60 0.94 1.02
Table 2: Comparison of PSNR, SSIM, and EPE between our method and Yan et al. (Yan et al. 2023) on RS-Homo.
RS-Homo dataset. The results in Table 2 show our diffusion models-based framework substantially improves upon Yan et al.’s state-of-the-art single-image model. Demonstrating great proficiency in managing challenging scenarios, our proposed method surpasses specifically designed architectures, enhancing the capability of diffusion models to address rolling shutter (RS) effects effectively.
@@
Yan et al. 18.48 0.55 4.18
RS-Diffusion 22.02 0.69 2.12
Table 3: Comparison of PSNR, SSIM, and EPE between our method and Yan et al. (Yan et al. 2023) on RS-Real.
@@
RS-Real dataset. Our dataset, designed for realism in content, RS-motion, and label accuracy, comprises 40,000 training and 1,000 test pairs across diverse scenes. It includes RS and GS images along with GT flow. We benchmark against Yan et al. (Yan et al. 2023) by retraining their model using the default settings on our trainset. The results confirm that our approach consistently outperforms theirs in both photometric and motion measures as demonstrated in Table 3.
Figure 6: Comparison on RS-Real and RS-Homo dataset. Column 1 shows the input RS image and Column 2 shows the ground-truth GS image. Column 3 and 5 are results by Yan et al. (Yan et al. 2023) and our method with the alignment heatmaps that use darker shades to indicate greater similarity compared to ground-truth GS images.
Qualitative Comparison
@@
Our method is evaluated against current methods on different benchmarks. We follow the settings used by Yan et al. (Yan et al. 2023) to compare with existing methods (Rengarajan, Balaji, and Rajagopalan 2017; Rengarajan, Rajagopalan, and Aravind 2016; Purkait, Zach, and Leonardis 2017; Grundmann et al. 2012; Kandula, Kumar, and Rajagopalan 2020) in Fig. 5, and ours outperforms the others. In addition, due to limited open-source options, we compare with Yan et al. (Yan et al. 2023) method on the remaining datasets. We visualize RS and ground-truth GS images in the first two columns, then present Yan et al. results paired with alignment heatmaps that use darker shades to indicate greater similarity compared to GT images. We then display our own results alongside their corresponding alignment heatmaps, illustrating our approach’s consistency with benchmarks and precision in alignment. This visual comparison highlights the effectiveness of our methodology.
RS-Real and RS-Homo Datasets. In the first three rows of Fig. 6, we present the results on the RS-Real dataset, while the last row shows the results on the RS-Homo dataset. Our approach consistently outperforms the method of Yan et al. (Yan et al. 2023) across varied RS challenges, including right-skewed images, dynamic scenes with moving cars, complex motion patterns from quick device movement, and low-texture scenarios, proving its robustness and superior correction capabilities.
Ablation Studies
@@
We evaluate our framework design through experiments, starting with comparisons under diverse generative objectives, paired with analytical analyses. The effectiveness of the Patch-Attention Block is then scrutinized.
Predicting IGF vs. GS Image
@@
Table 4: Comparison between different generating objects. “image-to-image” indicates transforming RS images to GS. “image-to-motion” represents predicting IGFs from RS.
We contrast our “image-to-motion” pipeline with traditional “image-to-image” methods that use diffusion models to convert RS images to GS images at a fixed resolution of 256$\times$256, later upscaled to 600$\times$800 for visual metric comparison with GT GS images, as shown in Table 4. Our approach outperforms the conventional framework in the RS removal task for two main reasons: 1) It is not constrained by diffusion model resolutions, since we can upsample IGF to correct RS images at their original size; 2) While “image-to-image” models learn the joint probability distribution between RS and GS images, our method gains additional improvements by learning from the distribution involving IGFs, RS and GS images, thus leveraging more information for enhanced results.
Rolling-Shutter Patch Attention
As demonstrated in Table 5, our proposed Patch Attention module proves to be effective, as it not only aggregates features within individual patches but also facilitates interaction between features across different patches. The experimental results align well with our prior understanding of RS motion.
@@
$\checkmark$ $\checkmark$ 22.02 0.69 2.12
Table 5: The effectiveness of Patch Attention. Intra refers to intra-patch attention and Inter indicates inter-patch attention.
Conclusion
In this work, we have presented RS-Diffusion, the first Diffusion Model based approach for single frame rolling shutter rectification. We have captured a novel dataset, namely RS-Real, to accomplish this task. The RS-Real dataset contains captured RS images, and the corresponding ground-truth GS images can be created according to synchronized gyroscope data that recorded during the RS frame capturing, yielding RS-GS image pairs. Uniquely, this dataset fulfills both accuracy and authenticity requirements for RS research. In addition, we have presented RS-Diffusion for real-time RS correction using just one RS frame. We have achieved state-of-the-art performances when compared to previous single RS correction methods. We publicly share our code and dataset with the community at https://github.com/lhaippp/RS-Diffusion.
Acknowledgements
This work was supported in part by National Natural Science Foundation of China (NSFC) under Grant Nos. 62372091, 62071097 and in part by Sichuan Science and Technology Program under Grant Nos. 2023NSFSC0462, 2023NSFSC0458, 2023NSFSC1972.